The observed rotation of 2.0g of a compound in a 50ml of solution in a polarimeter tube 20cm long is +138°, what is the specific rotation of the compound?
1
Expert's answer
2019-05-22T07:03:46-0400

Specific rotation is equal to: "[\alpha]=\frac{\alpha}{l\cdot c}"

where "\alpha" is the measured rotation in degrees, l is the path length in decimeters, c is the concentration in g/mL. Thus,

"[\alpha]=\frac{\alpha}{l\cdot c}=\frac{\alpha \cdot V}{l\cdot m}=\frac{+138^{\circ} \cdot 50 mL}{2dm \cdot 2.0 g}= +1725^{\circ}"